# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

### v1

@@ Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. It builds on the assumption that different forgery source domains exhibit distinct style statistics. Previous methods typically require fully fine-tuning pre-trained networks, consuming substantial time and computational resources. In turn, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild. The code is available at https://github.com/ChenqiKONG/OSDFD.
@@ I Introduction
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has led to significant progress in face forgery techniques, producing increasingly sophisticated fake faces that are indistinguishable from the human eye. Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1]. The recent emergence of diffusion and other advanced large vision-language models has exacerbated these challenges [2]. Consequently, developing effective detection methods to combat digital face forgery attacks has become critically important [3].
@@
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [4], color components [5], and DCT maps [6] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [7], head pose artifacts [8], and face-warping features [9]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [10], Efficient Net [11], and Capsule Net [12]. Later approaches, including Face X-ray [13], SBI [14], DCL [15], and RECCE [15] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [16], numerous ViT-based detectors [17, 18, 19, 20] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance.
However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In practical scenarios, detection models often require fine-tuning newly acquired data to adapt to new domains. However, adapting pre-trained detection models to new data is computationally costly and impedes deployment in real-world settings, especially on mobile devices with limited computing resources.
@@
To address these two challenges, this paper introduces a generalized and parameter-efficient approach for Open-Set DeepFake Detection (OSDFD). Different forgery-style statistics may exhibit unique properties critical for face forgery detection. Fig. 1 shows that existing forgery detectors have a significantly higher false negative rate (FNR) than the false positive rate (FPR) in unforeseen forgery domains, indicating a tendency to misclassify fake faces as real when encountering new domains in uncontrolled environments. Grounded in this observation, we propose a forgery style mixture module that enhances the diversity of source forgery domains by randomly mixing the feature statistics of different forgery styles, thereby effectively improving the model’s generalizability.
@@
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. Fig. 2 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
@@
  • We introduce a CDC adapter and a LoRA module into pre-trained plain ViT models. This design successfully captures forgery-related features and enables parameter-efficient face forgery detection.
  • We carry out extensive experiments and demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
@@ II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [7], head pose inconsistency [8], face warping artifacts [9], and rPPG features [22]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [10, 23, 24, 6], EfficientNet-b4 [11, 25], and Capsule network [12], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [13] and SBI [14] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [24] mined frequency artifacts, while DCL [15] and CFM [25] employed contrastive learning schemes for more generalized Deepfake detection. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [17, 18, 19, 26, 21, 27, 28] have been developed, showing superior detection performance. ICT [17] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
@@ III-B Forgery-aware PEFT
where $h^{in}\in\mathbb{R}^{d\times d}$, $W_{q/k/v}^{down}\in\mathbb{R}^{d\times r}$, $W_{q/k/v}^{up}\in\mathbb{R}^{r\times d}$, and $r<<d$. We optimize the slim, decomposed low-rank matrices during training instead of the redundant pre-trained weights. As such, the number of trainable parameters is significantly reduced. In this work, we set the hidden state dimension $d$ of ViT-B to 768, and the rank $r$ is defaulted to 8.
Figure 6: Pipeline of the proposed forgery feature mixture process.
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) they effectively preserve knowledge learned from ImageNet, while adaptively learning forgery-specific information; and (2) they simultaneously capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance. These advantages jointly contribute to a more generalized approach to face forgery detection.
@@ III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ denotes the source real faces, and $F_{i}^{S}$ is the $i_{th}$ source forgery domain, $i\in\{1,2,...,N\}$. As such, the testing data can be denoted as $D^{T}=\{R^{T},F_{1}^{T},F_{2}^{T},...,F_{M}^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains, while $M$ denotes the number of target forgery domains. Our preliminary experiment, illustrated in Fig. 1, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we need to augment source forgery domains to mitigate this issue. Fig. 5 depicts the proposed forgery-style mixture pipeline. The training batch $X^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$ is processed by our designed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are combined in the designed forgery-style mixture module, and the mixed features are forwarded to the classifier to perform forgery detection.
Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig. 6. This module processes a training batch of real and forged face images sampled from $N$ source domains, where $N=3$, for instance, is shown in Fig. 6. Green squares denote real faces in this representation, while red circles indicate the forgery faces. The input feature $f^{in}$ proceeds to our designed Forgery Style Mixture module.
We first sort $f^{in}$ to $f^{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from distinct forgery domains. The superscript of each sample indicates the feature index of the domain. To augment the forgery style, we first shuffle the fake features of $f^{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{f}^{sort}$ differs from the original $f^{sort}$. In addition, we keep the real feature order unchanged. Subsequently, we blend the forgery feature styles of the original $f^{sort}$ and the shuffled $\widetilde{f}^{sort}$.
@@
Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $f$ and $\tilde{f}$:
$\gamma_{mix}=\delta\sigma(f)+(1-\delta)\sigma(\tilde{f}),$ (7)
$\eta_{mix}=\delta\mu(f)+(1-\delta)\mu(\tilde{f}),$ (8)
where $\delta$ is a weight sampled from Beta distribution, i.e., $\delta\sim Beta(0.1,0.1)$. $\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $f$ and $\tilde{f}$. Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $f$:
$f_{mix}={\rm MixStyle}(f)=\gamma_{mix}\frac{f-\mu(f)}{\sigma(f)}+\eta_{mix}.$ (9)
Similar to [40], we set a default probability of 0.5 to activate the forgery-style mixture module. Finally, we restore the original index order of the mixed features $f^{sort}$ to obtain $f^{out}$, which is forwarded to the classifier for final decision-making. In the inference stage, the forgery-style mixture is deactivated.
During training, the designed forgery-style mixture module augments the source forgery domains, effectively mitigating the overfitting problem and enhance the model’s generalizability. This design enhances the model’s generalization for open-set face forgery detection.
@@ IV-A1 Implementation details
In this study, we employ the popular dlib face detector [47] to crop face regions. The proposed framework is implemented on the PyTorch [48] platform, and the model is trained using Adam optimizer [49] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. Our default architecture employs the ViT-B as the backbone, and we train the model on a single 2080Ti GPU with a batch size of 48. The training comprises 30,000 iterations, and the ViT-B parameters are initialized with ImageNet-21k [50] weights.

### v2

@@ Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. Our method builds on the assumption that different forgery source domains exhibit distinct style statistics. Specifically, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild.
@@ I Introduction
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has led to significant progress in face forgery techniques, producing increasingly sophisticated fake faces that are indistinguishable from the human eye. Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1, 2]. The recent emergence of diffusion and other advanced large vision-language models has exacerbated these challenges [3]. Consequently, developing effective detection methods to combat digital face forgery attacks has become critically important [4].
@@
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [5], color components [6], and DCT maps [7] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [8], head pose artifacts [9], and face-warping features [10]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [11], Efficient Net [12], and Capsule Net [13]. Later approaches, including Face X-ray [14], SBI [15], DCL [16], and RECCE [16] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance. However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In real-world applications, detection models typically need fine-tuning with newly acquired data to adapt to different domains. However, this adaptation process for pre-trained detection models can be computationally expensive, making it challenging to deploy them effectively, particularly on mobile devices with limited computational resources.
Figure 2: Prediction score distributions of real and fake faces of four unseen Deepfake datasets: DFDC, DFR, WDF, and FFIW. The three rows represent three Deepfake detectors: Xception, Efficient Net, and ViT-B. It can be observed that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces.
@@
To address these two challenges, this paper introduces a generalized and parameter-efficient approach for Open-Set DeepFake Detection (OSDFD). Different forgery-style statistics may exhibit unique properties critical for face forgery detection. Fig. 1 shows that existing forgery detectors have a significantly higher false negative rate (FNR) than the false positive rate (FPR) in unforeseen forgery domains, indicating a tendency to misclassify fake faces as real when encountering new domains in uncontrolled environments.
In Fig. 2, we further demonstrate that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces. Based on this observation, we propose a forgery style mixture module to increase the diversity of source forgery domains during training, thereby effectively mitigating forgery domain gaps in cross-dataset evaluations. Specifically, Fig. 2 presents cross-dataset experimental results for three representative face forgery detectors: Xception [11], EfficientNet [12], and ViT-B [17]. These models are trained on FF++ [22] and evaluated on DFDC [23], DFR [24], WDF [25], and FFIW [26]. In Fig. 2, green and red indicate real and fake faces, corresponding to ground-truth labels of ‘0’ and ‘1’, respectively. The $x$-axis and $y$-axis represent model prediction scores and the number of corresponding faces. As shown in Fig. 2, while these detectors perform well in identifying real faces, they struggle with fake faces in cross-dataset evaluations. This finding suggests that discrepancies between source and target datasets primarily arises from forgery domain gaps. Motivated by this observation, we design the forgery style mixture module to enhance the diversity of source forgery domains, thereby improving the model’s generalizability.
@@
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a notable contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation. Fig. 3 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
@@
  • Our approach explicitly captures forgery-related artifacts by integrating a CDC adapter and a LoRA module into a ViT-based detection framework. This targeted design enhances the model’s ability to extract both global and local forgery cues, which is crucial for Deepfake detection.
  • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection. Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
@@ II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [8], head pose inconsistency [9], face warping artifacts [10], and rPPG features [28]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [11, 29, 30, 7, 31], EfficientNet-b4 [12, 32, 33, 34], and Capsule network [13], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [14] and SBI [15] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [30] mined frequency artifacts, while DCL [16] and CFM [32] employed contrastive learning schemes for more generalized Deepfake detection. NoiseDF [35] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] have been developed, showing superior detection performance. ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
@@ III-B Forgery-aware PEFT
Figure 7: Pipeline of the proposed forgery feature mixture process.
@@
where $h^{in}\in\mathbb{R}^{d\times d}$, $W_{q/k/v}^{down}\in\mathbb{R}^{d\times r}$, $W_{q/k/v}^{up}\in\mathbb{R}^{r\times d}$, and $r<<d$. We optimize the slim, decomposed low-rank matrices during training instead of the redundant pre-trained weights. As such, the number of trainable parameters is significantly reduced. In this work, we set the hidden state dimension $d$ of ViT-B to 768, and the rank $r$ is defaulted to 8. The designed LoRA and CDC adapter respectively capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance.
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) By optimizing a small subset of parameters, our method reduces computational and memory requirements compared to full fine-tuning while also accelerating training, making frequent updates more practical in open-set environments. Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms. (2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various backbones in a plug-and-play fashion. (3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
@@ III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ represents the real source faces, while $F_{i}^{S}$ denotes the $i_{th}$ source forgery domain, where $i\in\{1,2,...,N\}$. Following previous methods in face forgery detection, we use FF++ as the training set, which includes multiple training forgery domains (i.e., manipulation types). As such, the target testing domain can be denoted as $D^{T}=\{R^{T},F^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains. Our preliminary experiments, illustrated in Fig. 1 and Fig. 2, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we propose to augment source forgery domains to mitigate this issue. Fig. 6 depicts the proposed forgery-style mixture pipeline. Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are passed to the designed forgery-style mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig. 7. Green squares denote real faces in this representation, while red circles indicate the forgery faces. This module processes each training batch consisting of real and fake face images sampled from the real source domain and $N$ forgery source domains. In Fig. 7, the subscripts of $f$ represent the domain labels, while the superscripts denote the sample index within each domain. We illustrate the case of NN=3 forgery domains in Fig. 7, with 3 samples per domain. The features $\mathcal{F}_{in}$ of each input batch is proceeded to our designed Forgery Style Mixture module.
We first sort $\mathcal{F}_{in}$ to $\mathcal{F}_{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from different forgery domains. To augment the forgery style, we first shuffle the fake features of $\mathcal{F}_{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{\mathcal{F}}_{sort}$ differs from the original $\mathcal{F}_{sort}$. In addition, we keep the real feature order unchanged. Subsequently, we mix the original $\mathcal{F}_{sort}$ and the shuffled $\widetilde{\mathcal{F}}_{sort}$ by fusing their feature statistics. The details of this process are provided next.
@@
Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$:
[…]
where $\delta$ is a weight sampled from Beta distribution, i.e., $\delta\sim Beta(0.1,0.1)$. $\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$. Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $\mathcal{F}_{sort}$:
[…]
Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making. Similar to [57], we set a default probability of 0.5 to activate the forgery-style mixture module. In the inference stage, the forgery-style mixture is deactivated. As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability. This design effectively improves the model’s ability to generalize in open-set face forgery detection.
@@
DCL [16] AAAI22 19.35M 0.949 0.877 0.829 0.684 - - - - - -
Implicit ID [64] CVPR23 22.36M 0.950 0.882 0.844 0.694 - - - - - -
UCF [65] CVPR23 20.81M - 0.867 - 0.679 - - - - - -
DiffusionFake [66] NIPS24 19.34M - 0.882 - 0.702 - - - - - -
@@
DCL [16] AAAI22 19.35M 0.838 0.759 0.751 0.679 - - - - - -
Implicit ID [64] CVPR23 22.40M 0.846 0.769 0.748 0.680 - - - - - -
UCF [65] CVPR23 20.81M - 0.746 - 0.673 - - - - - -
DiffusionFake [66] NIPS24 19.34M - 0.751 - 0.713 - - - - - -
@@ IV-A1 Implementation details
In this study, we employ the popular dlib face detector [67]. The detected face regions are enlarged by a factor of 1.3 and subsequently resized to 224×224 pixels before being input into the model. The proposed framework is implemented on the PyTorch [68] platform, and the model is trained using Adam optimizer [69] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. Our default architecture employs the ViT-B as the backbone, and the ViT-B parameters are initialized with ImageNet-21k [70] weights. We set the batch size to 48, as a single 2080Ti GPU can accommodate a maximum of 48 images at a resolution of $\times224$. During training, we use AUC as the validation metric. The training process consists of 30,000 iterations, as we observe that the validation AUC converges after about 20,000 iterations.